OpenStudy (anonymous):

The integral of cos^5 tdt Help!?!?

OpenStudy (jamesj):

First write \[ \cos^5 = \cos (1 - \sin^2)^2 \] Now you can substitute \( u = \sin x \) and it becomes fairly easy

OpenStudy (anonymous):

Can you solve it..Because Im totally black. :/

zepdrix (zepdrix):

\[\large \int\limits\limits \cos^5t \;dt \qquad = \qquad \int\limits \cos t \;\cos^2t \;\cos^2 t \;dt\]Using this identity,\[\large \color{salmon}{\cos^2t+\sin^2t=1}\qquad\rightarrow\qquad \color{salmon}{\cos^2t=1-\sin^2t}\] \[\int\limits\limits \cos t \;\color{salmon}{\cos^2t \;\cos^2 t} \;dt \qquad \rightarrow \qquad \int\limits \cos t\;\color{salmon}{(1-\sin^2t)(1-\sin^2t)}dt\] Which as James showed, we can write as,\[\large \int\limits \cos t(1-\sin^2t)^2 dt\]

zepdrix (zepdrix):

Let,\[\large \color{orangered}{u=\sin t}\]Taking the derivative gives us,\[\large \color{royalblue}{du=\cos t \;dt}\] Now plug those pieces into here to make the substitution. \[\large \int\limits\limits (1-\color{orangered}{\sin}^2\color{orangered}{t})^2 \color{royalblue}{\cos t\;dt}\]

zepdrix (zepdrix):

Giving us,\[\large \int\limits (1-\color{orangered}{u}^2)^2 \; \color{royalblue}{du}\]

zepdrix (zepdrix):

We'll expand out the brackets.\[\large \int\limits 1-2u^2+u^4 \; du\] Integrating, applying the Power Rule for Integration to each term, gives us,\[\large u-\frac{2}{3}u^3+\frac{1}{5}u^5+C\] Our final step will be to back-substitute. We need it back in terms of `t` since that's what we started with.

zepdrix (zepdrix):

\[\large \color{orangered}{u}-\frac{2}{3}\color{orangered}{u}^3+\frac{1}{5}\color{orangered}{u}^5+C \qquad \rightarrow \qquad \color{orangered}{\sin t}-\frac{2}{3}\color{orangered}{\sin}^3\color{orangered}{t}+\frac{1}{5}\color{orangered}{\sin}^5\color{orangered}{t}+C\]

zepdrix (zepdrix):

So there ya go. Lemme know if you're confused on any of those steps.